OpenStudy (kearsten):

gcf of 400x^2 and pi(64*pi*x^2)^2

OpenStudy (welshfella):

first look at the 2 numbers 400 and 64^2

OpenStudy (welshfella):

what is the greatest common factor of these 2 writting them as prime factors will help 400 = 2*2*2*2*5*5 64^2 = 2^12

OpenStudy (kearsten):

16?

OpenStudy (welshfella):

right

OpenStudy (welshfella):

now you just have the terms in x to deal with what is the GCF of x^2 and (x^2)^2?

OpenStudy (welshfella):

that is an easy one - it is of course x^2

OpenStudy (welshfella):

so our answer is 16x^2

OpenStudy (kearsten):

thanks